lexxii:

In parallelogram DEFG, DH = x + 3, HF = 3y, GH = 2x – 5, and HE = 5y + 2. Find the values of x and y. here is what I got x=3 and y=2 but those arent any of the options. What did I do wrong?

lexxii:

@Shadow

lexxii:

@theDeviliscoming

563blackghost:

Remember that opposite lengths are equal to each other. Where did the H come from though? The parallelogram is stated as `DEFG`

lexxii:

I forgot to add this sorry!

563blackghost:

Okay great a visual!

563blackghost:

I got the same thing >.> hrmmm

lexxii:

And the options are x = 36, y = 13 x = 39, y = 14 x = 14, y = 39 x = 13, y = 36 so I'm a bit confused

Shadow:

\[DH = HF\] \[x + 3 = 3y\] \[x = 3y - 3\] \[GH = HE\] \[2x - 5 = 5y + 2\] \[2(3y - 3) -5 = 5y + 2\] \[y = 13\] \[x + 3 = 3(13)\] \[x = 36\]

Shadow:

They bisect each other so each side is equal to each other.

lexxii:

Oh okay! Thank you so much!

563blackghost:

Oh okay, so they are in fact equal but I had to side it to one side to solve by substitution...i see e.e